what is vant hoff factor of salt al2(so4)3.........................

Grade:12

2 Answers

Umakant biswal
7 years ago
@ sandeep 
the compound will dissociate as 2 al3+  and 3so4^2- 
so, total no of ions after dissociation = 2+3 =5 
vant hoff factor = total no of ions after dissociation / before dissociation = 5/1 = 5 
so, here the vant hoff factor will be 5 . 
HOPE IT CLEARS YOUR DOUBT 
ALL THE BEST ..
Dhruvi Vakharia
4 years ago
Al2(SO4)3   =>   2Al3+ + 3SO42-
      1                  0          0         (initial)
     0                 2          3         (final after dissociation)
 
Van’t Hoff factor (i) = Moles after dissociaton/ Total initial moles
That is, (i) = (2+3)/1
Thus, (i)= 5